OpenStudy (anonymous):

Can someone check my double angle calculations please? The equation is below

OpenStudy (anonymous):

\[\sin 2\theta=\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

my answer is \[\sin \theta=\frac{ 1 }{ \sqrt{5} }\]

OpenStudy (unklerhaukus):

aren't you meant to find theta?

OpenStudy (anonymous):

well yes, but the question asked for Sin theta which is y/r :)

OpenStudy (anonymous):

Looks good assuming that theta is in the is in the first quad. actually it could be positive or negative.

OpenStudy (anonymous):

Use the half angle identity.

OpenStudy (anonymous):

@Paynesdad i did, but i'm unsure of my work.

OpenStudy (anonymous):

It is correct just include positive or negative.

OpenStudy (anonymous):

Alright! thanks! :)

OpenStudy (anonymous):

@Paynesdad Oh i forgot to mention why my answer was positive.. They told me that 2theta is between pi/2 (90degrees) and Pi (180 degrees) that excludes the negative answer correct?

OpenStudy (anonymous):

Yes because no theta is between pi/4 and pi/2 and thus in the first quadrant. Good work.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

No problem.